OpenStudy (anonymous):

Need someone to check my answer \[\int\limits\limits_{16}^{49}\frac{ \ \sqrt{lnx} dx }{ x } \]

OpenStudy (anonymous):

I got approximately 2.04

OpenStudy (anonymous):

would it help if I did the steps?

OpenStudy (badhi):

yes please do

OpenStudy (anonymous):

well i used substitution. \[u=\ln x\]\[du=\frac{ dx }{ x }\]

OpenStudy (anonymous):

\[\int\limits_{\ln16}^{\ln49}\sqrt{u}\times du\]

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }u ^{3/2}\]

OpenStudy (anonymous):

from ln 16 to ln 49

OpenStudy (anonymous):

and i got 2.04

OpenStudy (anonymous):

You did well then.

OpenStudy (anonymous):

does this seem correct?

OpenStudy (anonymous):

THe limits are 49 and 16 tho, not ln 49 and ln 16

OpenStudy (anonymous):

he did an u substitution @Meepi, hence his limits are correct: \[\Large u=\ln x \] from x=16 to x=49

OpenStudy (anonymous):

Your answer is correct @MATTW20

OpenStudy (anonymous):

yeah wasn't looking my bad

OpenStudy (anonymous):

Thank You :)

OpenStudy (anonymous):

you're welcome